Which equation is equivalent to y=7x^5?
A. Ln y= ln 7 + 5 ln x
B. Y= ln 7 + 5 ln x
C. Y= 7 + 5 ln x
D. Ln y = 7 + 5 ln x
E. Ln y = ln 7 - 5 ln x

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Which equation is equivalent to y=7x^5?
A. Ln y= ln 7 + 5 ln x
B. Y= ln 7 + 5 ln x
C. Y= 7 + 5 ln x
D. Ln y = 7 + 5 ln x
E. Ln y = ln 7 - 5 ln x
==============================
If y = a*x^b
ln(y) = ln(a) + b*ln(x)
===========
PS Use parentheses.

Answer by ikleyn(52781)

.
The given equation and the equations A) - E) have different DOMAINS.

Therefore, the given equation IS NOT equivalent to any of equations A) - E).